# Bias in Decision-Making for AI’s Ethical Dilemmas: A Comparative Study of ChatGPT and Claude

### v1

@@ Introduction
$S_{\alpha,\beta}=1-\frac{\sum_{pa_{1}\in G_{\alpha},pa_{2}\in G_{\beta}}N_{pa_{1},pa_{2}}}{50}$ (5)
, where $S_{\alpha,\beta}$ is the unselected frequency of the protected attribute group $G_{\alpha,\beta}$, 50 is the number of times we asked LLMs in one round. For example, the protected attributes in the group $G_{Gender,Color}$ were selected 45 times, and the unselected frequency of the protected attribute group $G_{Gender,Color}$ is 0.1.
Then we calculated the normalized unselected frequency of the specific single protected attribute group using:
$S^{*}_{\gamma}=\frac{\sum_{\alpha=\gamma}S_{\alpha,\beta}}{Count_{\alpha=\gamma}}$ (6)
, where $S^{*}_{\gamma}$ is the normalized unselected frequency of the specific single protected attribute group $G_{\gamma}$ for intersectional scenario, $Count_{\alpha=\gamma}$ is the number of the intersectional protected attribute groups including the specific single protected attribute group. For example, $S^{*}_{Gender}$ is the normalized unselected frequency of the specific single protected attribute group $G_{Gender}$ for intersectional scenario, $Count_{\alpha=Gender}$ is the number of the intersectional protected attribute groups including the specific single protected attribute group and is 6, and $\sum_{\alpha=\gamma}S_{\alpha,\beta}$ is 1.2, and the normalized unselected frequency of the specific single protected attribute group $G_{Gender}$ for intersectional scenario is 0.2.
By calculating the results of the 4 rounds of experiments, we got the mean unselected frequency of each single protected attribute group for single protected attribute scenario and the mean normalized unselected frequency of each single protected attribute group for intersectional scenario.
@@
$\sigma_{\alpha,\beta}=\sum_{pa_{1},pa_{2}\in G_{\alpha,\beta}}\sigma_{pa_{1},pa_{2}}$ (7)
, where $\sigma_{\alpha,\beta}$ is the total standard deviation of group $G_{\alpha,\beta}$, $\sigma_{pa_{1},pa_{2}}$ is the standard deviation of intersectional protected attribute $pa_{1},pa_{2}$.
@@
$\sigma^{*}_{\gamma}=\frac{\sum_{\alpha=\gamma}\sigma_{\alpha,\beta}}{Count_{\alpha=\gamma}}$ (8)
, where $\sigma^{*}_{\gamma}$ is the normalized total standard deviation of the specific single protected attribute group $\gamma$ for intersectional protected attribute scenario, $Count_{\alpha=\gamma}$ is the number of the intersectional protected attribute groups including the specific single protected attribute group $\gamma$.
@@
$\mu_{a}=\frac{\sum_{pa\in a}f^{*}_{pa}}{n_{a}}$ (9)
,where $\mu_{a}$ is the cluster’s mean value, and $f^{*}_{pa}$ is the mean normalized frequency of protected attribute $pa$, $n_{a}$ is the number of samples in cluster $a$.
@@
$\Delta SSE=\frac{n_{a}n_{b}}{n_{a}+n_{b}}(\mu_{a}-\mu_{b})^{2}$ (10)
, where $\Delta SSE$ is the increase in intra-cluster variance after merging two clusters, $n_{a}$ and $n_{b}$ are the number of samples in the two clusters, and $\mu_{a}$ and $\mu_{b}$ are the two clusters’ mean values.
(a) Standard deviation comparison for single protected attribute scenario
(b) Standard deviation comparison for intersectional protected attribute scenario
Figure 4: Stacked bar visualization comparing standard deviations between single protected attributes (panel a) and intersectional protected attributes (panel b). Protected attributes along the x-axis are arranged in ascending order based on the cumulative standard deviations of GPT-3.5 Turbo and Claude 3.5 Sonnet, with contributions from each model shown in distinct colors.
@@
(a) GPT clustering
(b) Claude clustering
Figure 5: Hierarchical clustering of GPT-3.5 Turbo (panel a) and Claude 3.5 Sonnet (panel b). Cluster 1 exhibits high mean normalized frequencies, while Cluster 2 demonstrates low mean normalized frequencies, and Cluster 3 displays the highest mean normalized frequency.
We calculated the preference score for each protected attribute. If the score is positive, GPT-3.5 Turbo prefers that protected attribute. Otherwise, Claude 3.5 Sonnet prefers the protected attribute. A larger absolute value of the score indicates that the LLM prefers the protected attribute over another LLM. For single protected attribute scenario, we calculated the preference score of the protected attribute using:
$B_{pa}=\frac{f_{pa}^{GPT}-f_{pa}^{Claude}}{f_{pa}^{GPT}+f_{pa}^{Claude}}$ (11)
, where $B_{pa}$ is the preference score of protected attribute $pa$, $f_{pa}^{GPT}$ and $f_{pa}^{Claude}$ are the mean normalized frequencies of protected attribute $pa$ for GPT-3.5 Turbo and Claude 3.5 Sonnet. For example, $B_{Masculine}$ is the preference score of protected attribute $Masculine$, $f_{Masculine}^{GPT}$ and $f_{Masculine}^{Claude}$ are the mean normalized frequencies of protected attribute $Masculine$ for GPT-3.5 Turbo and Claude 3.5 Sonnet and are 0.5 and 0.3, and the preference score $B_{Masculine}$ is 0.25.
@@ Results
Divergent Gender Biases and Demographic Preferences Across LLM Models
Figure 1 shows the mean normalized frequencies for GPT-3.5 Turbo and Claude 3.5 Sonnet. The x-axis and y-axis coordinates are labeled with 20 protected attributes from seven protected attribute groups. Each box represents the mean normalized frequency of an intersectional protected attribute consisting of the corresponding x-axis and y-axis coordinate protected attributes. The brighter the color, the higher the frequency; the darker the color, the lower the frequency. Since different sequential combinations of the same protected attributes are considered identical, all images are symmetric about the diagonal line. We then identified the top 10 brightest protected attributes to determine the 10 protected attributes that LLMs tend to favor the most.
Figure 1 and Figure 1 show the top 10 mean normalized frequency rankings for the intersectional protected attributes for both models, with the y-axis labeled as the intersectional protected attribute and the bar lengths representing the mean normalized frequency for that protected attribute. Figure 2 shows the mean normalized frequency rankings of single protected attributes for intersectional scenario, figure 2 for GPT-3.5 Turbo and figure 2 for Claude 3.5 Sonnet.
We identified the high-frequency protected attributes in Figure 1 and Figure 1 and sorted them to obtain Figure 1 and Figure 1, revealing the attributes favored by the LLMs. Furthermore, we calculated the selected frequency of single protected attributes and arranged these attributes according to their frequency in GPT-3.5 Turbo, from highest to lowest. Our findings indicate that: (1) The “Good-looking” appearance attribute shows high frequency in both models, while “Standard-looking” and “Unpleasant-looking” show extremely low frequencies. (2) The “Race” categories of “White”, “Black”, and “Yellow” in both models rank from high to low frequency, with “Yellow” showing particularly low frequency. (3) The “Age” categories of “8”, “35”, and “70” years in both models display decreasing frequency distribution. (4) For GPT-3.5 Turbo, the frequency of “Masculine” is notably high, while “Feminine” and “Androgynous” frequencies are very low; conversely, for Claude 3.5 Sonnet, “Masculine” frequency is very low, while “Feminine” and “Androgynous” frequencies are high.
Differential Sensitivity Patterns for Single and Intersectional Protected Attribute Scenarios
Figure 3 presents a detailed comparison of ethical sensitivity patterns between single and intersectional protected attribute scenarios for GPT-3.5 Turbo and Claude 3.5 Sonnet. For single protected attribute scenario (Figure 3), “Race” and “Color” emerge as highly sensitive attributes, with combined sensitivity scores reaching approximately 0.85 and 0.75 respectively. This heightened sensitivity suggests both models are particularly cautious when making ethical decisions involving these attributes in isolation. Other attributes such as “Disability”, “Gender”, “Dressing”, “Look”, and “Age” demonstrate notably lower sensitivity levels (below 0.2), indicating the models are more willing to make distinctions based on these attributes. However, the pattern shifts dramatically for intersectional protected attribute scenario (Figure 3), where sensitivity levels become more uniform across all protected attributes, ranging between 0.2 and 0.4. This equalization effect is notable for “Race” and “Color”, which show substantially reduced sensitivity when combined with other attributes, while previously less sensitive attributes like “Age” and “Disability” show increased sensitivity in intersectional contexts. This transformation in sensitivity patterns between single and intersectional scenarios suggests that the models’ ethical decision-making processes become more nuanced and balanced when considering multiple protected attributes simultaneously.
(a) Preference score for single protected attribute scenario
(b) Preference score for intersectional protected attribute scenario
Figure 6: Comparative preference score distributions for single (panel a) and intersectional (panel b) protected attribute scenario across GPT-3.5 Turbo and Claude 3.5 Sonnet. The preference scores, scaled from $-1$ to $1$, quantify algorithmic response biases between the two language models, where positive values (up to 1) indicate stronger preferences by GPT-3.5 Turbo, negative values (down to -1) represent stronger preferences by Claude 3.5 Sonnet, and zero indicates balanced preference between the models.
Stability Analysis of Protected Attribute Responses
Figure 4 presents a comparative analysis of model stability through total standard deviation histograms across single (Figure 4) and intersectional (Figure 4) protected attribute scenarios. The x-axis displays seven protected attribute groups (Age, Disability, Dressing, Race, Look, Gender, and Color), with stacked bars representing the standard deviations for GPT-3.5 Turbo (blue) and Claude 3.5 Sonnet (beige). For single protected attribute scenario (Figure 4), we observe varying degrees of stability across different attributes, with “Color” showing the highest combined standard deviation (approximately 0.8) and “Age” demonstrating the lowest (approximately 0.2). The intersectional scenario (Figure 4) reveals more consistent standard deviations across all attributes, with values generally ranging between 0.3 and 0.4, indicating more uniform stability. The calculation methodology for these standard deviations is detailed in the Ethical stability subsection of Measurement Metrics. Notably, the intersectional scenarios demonstrate more consistent stability patterns across all protected attributes compared to single scenarios, suggesting that the models’ responses are more predictable when dealing with combinations of protected attributes rather than individual ones. This enhanced stability in intersectional scenarios provides strong evidence for the reliability of our results.
Hierarchical Clustering Analysis of Model Preferences
Figure 5 illustrates the clustering hierarchical tree diagrams comparing GPT-3.5 Turbo (Figure 5) and Claude 3.5 Sonnet (Figure 5) for intersectional protected attribute scenario. The hierarchical clustering methodology is detailed in the Clustering of preference subsection of Measurement Metrics. The dendrograms visualize the clustering of protected attributes based on their preference strengths, with the y-axis representing the distance measure (from 0 to 2.5) indicating the degree of similarity between clusters. The analysis employs a 3-cluster classification system, wherein the ethical preferences of protected attributes within the same category demonstrate high proximity. Both models show distinct clustering patterns: GPT-3.5 Turbo exhibits a polarized distribution with a weak preference cluster containing appearance-related attributes and demographic characteristics (“70”, “Yellow”, “African”), a strong preference cluster encompassing most demographic attributes (“Disabled”, “Asian”, “Caucasian”), and a very strong preference exclusively for “Good-looking”. In contrast, Claude 3.5 Sonnet demonstrates a more balanced distribution, with a dominant strong preference cluster including most protected attributes (“Disabled”, “Modest”, “Feminine”), while maintaining the very strong preference for “Good-looking” and relegating attributes such as “Masculine” and “Luxury” to the weak preference cluster. This distinct clustering pattern suggests potentially different underlying biases in their respective training processes.
Comparative Analysis of Model Preferences
Figure 6 illustrates the preference score distributions for GPT-3.5 Turbo and Claude 3.5 Sonnet across single and intersectional scenarios. The preference scores, ranging from -1 to 1, quantify the relative biases between the two models, where a score of 1 indicates exclusive preference by GPT, -1 indicates exclusive preference by Claude, and 0 represents equal preference between models. For single protected attribute scenario (Figure 6), GPT demonstrates strong preferences (positive scores) for several attributes, with “Disabled”, “Black”, and “Caucasian” showing the highest positive scores ($>0.5$). Conversely, Claude shows distinct preferences (negative scores) for attributes including “Unpleasant-looking”, “Masculine”, and “Non-disabled”. This pattern suggests fundamental differences in the models’ ethical decision-making processes for individual protected attributes. The intersectional scenario (Figure 6) reveals interesting shifts in these preferences. “Masculine” and “African” attributes maintain strong positive scores for GPT, while “Feminine” and “Androgynous” remain strongly preferred by Claude. However, only 7 protected attributes (4 for GPT, 3 for Claude) maintain their preference direction across both scenarios, while 12 attributes (7 for GPT, 5 for Claude) show reversed preferences between single and intersectional contexts. This preference instability suggests that the models’ ethical decision-making processes become more complex and potentially less consistent when evaluating multiple protected attributes simultaneously. The visualization demonstrates the dynamic nature of these preferences through bar lengths, which represent the magnitude of preference disparity between the models. The clear separation of preferences in both scenarios, particularly for certain demographic and appearance-based attributes, highlights systematic differences in how these models approach ethical decision-making tasks.
@@
This study investigated the ethical decision-making patterns for GPT-3.5 Turbo and Claude 3.5 Sonnet through simulated ethical dilemmas, evaluating their ethical preferences, sensitivity, stability, and preference clustering. Our analysis revealed inherent biases in both LLMs’ decision-making processes, paradoxically making them more “human-like” but raising significant ethical concerns, particularly regarding their impact on underrepresented groups.
A striking finding across both models is the dominant role of physical appearance in ethical decision-making. Both LLMs showed strong preferences for “Good-looking” individuals, with this attribute consistently ranking high for both single and intersectional protected attribute scenarios. Moreover, each model exhibited distinct ethical decision-making patterns respectively. GPT-3.5 Turbo showed a clear preference hierarchy favoring stereotypically dominant groups (Non-disabled, Masculine, Middle-aged, White/Caucasian men). In contrast, Claude 3.5 Sonnet demonstrated more balanced preferences across diverse protected attributes including “Race”, “Age”, “Gender”, and “Disability” status. These preferences are irrelevant to ethical decision-making in driving-context scenarios.
Our comparative analysis of single versus intersectional scenarios revealed that ethical sensitivity significantly decreases in more complex scenarios. This finding extends beyond existing literature, which typically focuses on single protected attributes, highlighting the importance of examining intersectional biases in LLMs. This complexity could potentially be exploited to manipulate LLMs’ ethical decisions through feature combinations.
A notable discovery concerns the impact of linguistic referents on ethical evaluations. Our analysis revealed that terminology choices significantly influence model preferences – for instance, ”Asian” received higher preference scores than “Yellow” in ethical dilemmas. This aligns with research on covert racism in LLMs and suggests that linguistic choices play a crucial role in shaping model biases (Hofmann et al. 2024). The lower preference for color-based descriptors may reflect historical discriminatory practices and ongoing underrepresentation of certain communities.
@@
These biases raise concerns about the integration of LLMs into autonomous systems (e.g., self-driving vehicles, autonomous weapons, disaster response systems). Such biases may stem from training data imbalances reflecting real-world underrepresentation. While the precise mechanisms behind these biases remain unclear due to limited model transparency, our findings emphasize the moral imperative for human oversight in LLM development and deployment.
Future Work
Future research directions emerge in three key areas. First, expanding the scope to include a broader range of protected attributes, LLM types (including open-source models), and linguistic contexts could enhance our understanding of bias patterns across different societal and cultural backgrounds. Second, investigating the impact of prompt engineering on ethical decision-making could reveal how variations in tone, phrasing, and emotional content influence model responses. Finally, comparative studies between human and LLM ethical decision-making could illuminate the extent of moral alignment between artificial and human intelligence, providing crucial insights for advancing human-AI interaction in ethical contexts.

### v3

@@ Introduction
$S_{\alpha,\beta}=1-\frac{\sum_{pa_{1}\in G_{\alpha},pa_{2}\in G_{\beta}}N_{pa_{1},pa_{2}}}{10},$ (5)
where $S_{\alpha,\beta}$ is the unselected frequency of the protected attribute group $G_{\alpha,\beta}$, 10 is the number of times we asked LLMs in one round. For example, the protected attributes in the group $G_{Gender,Color}$ were selected 9 times, and the unselected frequency of the protected attribute group $G_{Gender,Color}$ is 0.1.
Then we calculated the normalized unselected frequency of the specific single attribute group using:
$S^{*}_{\gamma}=\frac{\sum_{\alpha=\gamma}S_{\alpha,\beta}}{Count_{\alpha=\gamma}},$ (6)
where $S^{*}_{\gamma}$ is the normalized unselected frequency of the specific single protected attribute group $G_{\gamma}$ for intersectional scenario, $Count_{\alpha=\gamma}$ is the number of the intersectional protected attribute groups including the specific single protected attribute group.
By calculating the results of the five rounds of experiments, we got the mean unselected frequency of each single protected attribute group for single protected attribute scenario and the mean normalized unselected frequency of each single protected attribute group for intersectional scenario.
@@
(a) Single attribute frequency heat map
(b) Intersectional attribute frequency heat map
Figure 1: Comparative heat maps of mean normalized frequency (0-1 scale) for single (a) and intersectional (b) protected attributes. Higher values indicate greater selection frequency, representing preference in protective dilemmas but rejection in harmful scenarios.
@@
$\sigma_{\alpha,\beta}=\sum_{pa_{1},pa_{2}\in G_{\alpha,\beta}}\sigma_{pa_{1},pa_{2}},$ (7)
where $\sigma_{\alpha,\beta}$ is the total standard deviation of group $G_{\alpha,\beta}$, $\sigma_{pa_{1},pa_{2}}$ is the standard deviation of intersectional protected attribute $pa_{1},pa_{2}$.
@@
$\sigma^{*}_{\gamma}=\frac{\sum_{\alpha=\gamma}\sigma_{\alpha,\beta}}{Count_{\alpha=\gamma}},$ (8)
where $\sigma^{*}_{\gamma}$ is the normalized total standard deviation of the specific single protected attribute group $\gamma$ for intersectional protected attribute scenario, $Count_{\alpha=\gamma}$ is the number of the intersectional protected attribute groups including the specific single protected attribute group $\gamma$.
@@
$\mu_{a}=\frac{\sum_{pa\in a}f^{*}_{pa}}{n_{a}},$ (9)
where $\mu_{a}$ is the cluster’s mean value, and $f^{*}_{pa}$ is the mean normalized frequency of protected attribute $pa$, $n_{a}$ is the number of samples in cluster $a$.
@@
$\Delta SSE=\frac{n_{a}n_{b}}{n_{a}+n_{b}}(\mu_{a}-\mu_{b})^{2},$ (10)
where $\Delta SSE$ is the increase in intra-cluster variance after merging two clusters, $n_{a}$ and $n_{b}$ are the number of samples in the two clusters, and $\mu_{a}$ and $\mu_{b}$ are the two clusters’ mean values.
@@
We calculated the preference score for each protected attribute. If the score is positive, Open-source LLMs prefer that protected attribute. Otherwise, Closed-source LLMs prefer the protected attribute. A larger absolute value of the score indicates that LLMs of one type prefer the protected attribute over LLMs of another type. For single protected attribute scenario, we calculated the preference score of the protected attribute using:
$B_{pa}=\frac{f_{pa}^{Open}-f_{pa}^{Closed}}{f_{pa}^{Open}+f_{pa}^{Closed}},$ (11)
where $B_{pa}\in[-1,1]$ is the preference score of protected attribute $pa$, $f_{pa}^{Open}$ and $f_{pa}^{Closed}$ are the mean normalized frequencies of protected attribute $pa$ for Open-source LLMs and Closed-source LLMs. For example, $B_{Black}$ is the preference score of protected attribute $Black$, $f_{Black}^{Open}$ and $f_{Black}^{Closed}$ are the mean normalized frequencies of protected attribute $Black$ for Open-source LLMs and Closed-source LLMs and are 0.5 and 0.3, and the preference score $B_{Black}$ is 0.25.
@@
(a) Sensitivity for single scenarios
(b) Sensitivity for intersectional scenarios
Figure 2: Comparative sensitivity analysis using heat maps for single (a) and intersectional (b) protected attributes.
(a) Standard deviation heat map for single scenarios
(b) Standard deviation heat map for intersectional scenarios
Figure 3: Comparative heat maps of standard deviations in single (a) and intersectional (b) attribute scenarios.
(a) Protective closed
(b) Protective open
(c) Harmful closed
(d) Harmful open
Figure 4: Hierarchical clustering of closed-source (a, c) and open-source LLMs (b, d) in protective/harmful dilemmas, revealing three distinct clusters by normalized frequency (Cluster 1: low, Cluster 2: high, Cluster 3: highest).
(a) Preference score for protective dilemmas
(b) Preference score for harmful dilemmas
Figure 5: We compare the preference scores (-1 to 1) of intersectional attributes between open-source and closed-source LLMs in (a) protective and (b) harmful dilemmas. Positive values indicate closed-source preference, negative values open-source preference, and zero denotes balanced responses.
@@ Results
What Are the General Tendency Characteristics of LLMs?
Figure 1 shows averaged heat maps of attribute selection frequency (0-1) for closed-source and open-source LLMs in four dilemmas. Higher values indicate stronger feature preference in protective scenarios but lower preference in harmful ones, revealing how different LLMs prioritize features across conditions.
Our analysis reveals that LLMs consistently exhibit distinct preferences, particularly in intersectional scenarios, favoring attributes like “Modest”, “Black”, “Good-looking”, and “Disabled” while avoiding features such as “Luxury”, “Yellow”, “Unpleasant-looking”, and “Non-disabled”. This pattern clearly demonstrates systematic biases in LLMs’ characteristic evaluations.
Figure 2 compares open-source and closed-source LLMs’ ethical sensitivity (on a 0-1 scale) across four scenarios. Both single (2) and intersectional (2) analyses show strongest sensitivity to “Gender”, “Color”, and “Race”, with weaker responses to “Age”, “Dressing”, and “Look”, revealing systematic prioritization biases in LLMs’ ethical frameworks.
Figure 3 illustrates the stability heatmaps of open-source and closed-source LLMs across four different scenarios, examining both single-feature (3) and intersectional-feature (3) cases. Larger values indicate poorer model stability. The results show that the standard deviation values of the models are generally low, which suggests that the propensity of the LLMs’ features is relatively stable. This observed stability lays the foundation for subsequent analyses.
Figure 4 displays clustering hierarchies of closed-source and open-source LLMs in protective (4, 4) and harmful (4, 4) dilemmas. In protective dilemmas, both model types show similar two-cluster structures: higher clusters contain preferred features (e.g., “Black”, “Good-looking”), while lower clusters include less preferred ones (e.g., “Unpleasant-looking”, “Yellow”). Harmful dilemmas reveal three clusters, with model-agnostic groupings: lower (“Feminine”, “Caucasian”), higher (“Non-disabled”, “Androgynous”), and highest (“Luxury”, “70”). These consistent patterns across model types suggest LLMs share fundamental value judgments, despite dilemma-dependent clustering differences.
What Are the Differences in Ethical Tendencies Between Open-Source and Closed-Source LLMs?
Closed-source LLMs reject “Androgynous” and prefer “Feminine” gender representations, while open-source LLMs show no strong gender preference (Figure 1). According to Figure 2, open-source LLMs demonstrate generally lower sensitivity than closed-source LLMs in protective scenarios. However, in harmful scenarios, open-source LLMs exhibit a significant sensitivity increase, compared to only marginal increases in closed-source models. This pattern suggests that open-source LLMs actively express ethical preferences by making deliberate choices in protective scenarios, while systematically avoiding harmful feature selections in harmful scenarios. Although both types of LLMs demonstrate good stability (Figure 3), open-source LLMs show relatively greater stability in single-factor scenarios, while both model types exhibit nearly equivalent stability in two-factor scenarios. Moreover, in Figure 5, we observe significant differences ($p<0.05$) in the tendencies of closed-source and open-source models across different scenarios. Open-source models exhibit stronger preferences for features such as “African”, “Disabled”, “Black”, and “Androgynous” compared to closed-source models. Meanwhile, closed-source models show weaker tendencies in protective scenarios but demonstrate distinct preferences for “Good-looking”, “Modest”, and “Masculine” in harmful scenarios. Our significance testing demonstrates that the differences between the two model types are statistically significant at the 0.05 level. These findings highlight the divergent behavioral patterns between the two types of models under varying conditions.
How Do Different Types of Dilemmas Affect Ethical Decisions of LLMs?
The propensity characteristics of LLMs remain consistent across protective scenarios but vary significantly in harmful scenarios for certain feature groups (Figure 1). For instance, in harmful scenario 1, LLMs predominantly target “Luxury” and “70”, whereas in harmful scenario 2, they primarily select “Modest” and “8”. This pattern suggests that while LLMs’ behavior remains consistent in protective contexts, their exclusion patterns in harmful scenarios differ according to specific scenario characteristics. From a sensitivity perspective, harmful scenarios elicit significantly higher sensitivity than protective scenarios (Figure 2), demonstrating that harmful contexts are more cognitively demanding for LLMs. This further suggests that LLMs find harmful decisions substantially more challenging than protective decisions. LLMs show comparable stability across different scenario types (Figure 3), indicating that this approach provides a reliable method for evaluating ethical tendencies of LLMs across diverse contexts.
How Does Feature Dimensionality Affect the Ethical Decision-Making of LLMs?
The propensity patterns of LLMs exhibit minimal variation between single-factor and two-factor scenarios (Figure 1). However, the two-factor condition more accurately reflects realistic behavioural tendencies, as LLMs demonstrate a capacity to conceal their tendencies when presented with simpler, single-factor choices (e.g., exhibiting restrained selection frequencies for attributes like “Good-looking” within the “Look” feature group). In contrast, the increased complexity of two-factor scenarios necessitates more sophisticated content processing, substantially impeding LLMs’ ability to regulate dispersed feature selection frequencies that might otherwise obscure intrinsic biases. This consequently reveals more distinct ethical tendencies (e.g., markedly stronger preferences for “Good-looking” in the “Look” feature group under two-factor conditions). These observations collectively indicate that near-authentic, complex feature interactions provide more reliable representations of the genuine behaviour propensities exhibited by LLMs. As shown in Figure 2, the sensitivity of LLMs in two-factor scenarios demonstrates greater homogeneity across different feature groups compared to single-factor scenarios. Specifically, feature groups exhibiting lower sensitivity in single-factor conditions show increased sensitivity in two-factor conditions, while those with higher sensitivity in single-factor conditions display reduced sensitivity in two-factor contexts. Analysis of Figure 3 reveals that while certain cases demonstrate low stability in single-factor scenarios, the two-factor experimental conditions exhibit more uniform and overall improved stability compared to single-factor conditions.
@@
This study investigated the LLMs’ ethical decision-making through simulated ethical dilemmas, evaluating their ethical preferences, sensitivity, stability, and preference clustering. Our analysis revealed inherent biases in all LLMs’ decision-making processes, which depend on both the type of dilemma and the model used. By doing so, this study raises significant ethical concerns when LLMs are applied in real-world decision-assistive or autonomous settings considering they may produce unfair and unstable recommendations to specific individuals or groups. By highlighting how preference patterns change across contexts, this work calls for public awareness of the potential risks in LLMs’ decision-making, urges comprehensive bias examination and audits, transparent and fair model development, and robust oversight to ensure generative AI makes fair and reliable decisions.
The foremost finding of this study is that all LLMs, whether open-source or closed-source exhibit consistent biases toward certain protected attributes, such as Dressing (“Modest”), Color (“Black”), Look (“Good-looking”), and Disability (“Disabled”). These preferences are persistent across multiple dilemma types even though these traits should not influence in ethical decisions like those in driving situations. This raises important concerns about the alignment of LLMs with societal fairness principles since they are irrelevant to ethical decision-making in the scenarios proposed in this study. For example, the strong bias toward “Good-looking” individuals highlights that these models may be influenced by superficial or culturally shaped notions, even when making serious ethical decisions.
Second, we observed notable differences between model types, not only in their preferences for specific features, but also in their levels of ethical sensitivity. Open-source models exhibited significantly stronger preferences for features such as African, Disabled, Black, and Androgynous compared to closed-source models, which displayed minimal bias in protective scenarios but pronounced tendencies toward “Good-looking”, “Modest”, and “Masculine” features in harmful scenarios. LLMs exhibit different level of ethical sensitivity depending on different types of ethical dilemmas. Such differences indicate a built-in tendency to avoid making morally harmful choices, perhaps due to their more diverse or regulated training corpora. This finding has practical implications for the deployment of LLMs in ethically sensitive domains. For instance, in automated systems where decisions may impact human safety (e.g., autonomous driving, emergency response), the choice between open-source and closed-source models could lead to different ethical outcomes. Attention must also be paid to the prevalent tendencies associated with each model type when deploying LLMs. The observed biases in closed-source models also raise concerns about their use in areas involving gender diversity and inclusion, such as education, healthcare, or customer service.
Additionally, the higher ethical stability of open-source LLMs in simpler conditions (single attribute) makes them potentially more reliable. However, their comparable performance under complex conditions (intersectional attributes) highlights the limitations of current LLMs in handling real-world complexity and supports the need for intersectionality-aware evaluation frameworks in AI ethics research.
Third, regarding to dilemma types, our findings show that LLMs respond more consistently in protective dilemmas but behave variably in harmful ones. In harmful scenarios, different features (e.g., Age, Dressing) are targeted depending on the situation, suggesting that LLMs’ ethical decisions are highly context-dependent when harm is involved. Additionally, LLMs show higher sensitivity in harmful dilemmas, indicating that such situations are more ethically challenging for them. This raises concerns for real-world use in critical moments involving human lives, like autonomous vehicles or medical rescue, where models must make quick, morally loaded decisions. If LLMs react inconsistently across harmful contexts, they may pose risks of unfair or biased outcomes. These results highlight the need to test LLMs across diverse dilemma types and improve their robustness in morally complex situations.
@@
The presence of bias in LLMs can lead to unfair, discriminatory, or even harmful outcomes in real life, especially when they are integrated into high-stakes scenarios such as healthcare and education, and even life-critical scenarios such as autonomous systems and disaster response systems. Such biases may stem from training data imbalances reflecting real-world underrepresentation. While the precise mechanisms behind these biases remain unclear due to limited model transparency, our findings emphasize the moral imperative for human oversight in LLMs’ development and deployment.
To address the risks of LLMs’ biases, this study also calls for greater awareness and action from involved stakeholders in LLMs development, deployment, and regulation.
Developers should carefully select training datasets and avoid algorithms that reinforce human stereotypes. Model training should incorporate fairness-aware techniques and balanced representation across protected groups. Besides, developers, especially of closed-source systems, should be more transparent about training data and fine-tuning methods to allow external auditing and understanding of model behavior in moral decision-making.
Policymakers should take ethical bias seriously when establishing standards for LLM deployment. For example, models that exhibit unacceptable ethical bias should not be allowed for direct use without essential safeguards. Regulatory frameworks should include bias testing, ensure transparency in model development and use, and provide reminders or disclaimers on specific occasions.
The public and users should remain cautious and critical of LLMs, especially in ethical decision-making contexts. Model outputs may reflect biased tendencies and should not be assumed to be objective or morally neutral. Raising public awareness of the limitations of AI is essential to prevent over-reliance and potential harm.
Limitations and Future Work
While this work contributes to the exploration of potential biases in several LLMs in simulated scenarios, it also has several limitations that require further work.
First, this study investigated several selected attributes in English-speaking context, which may limit the generalizability of the results. Future studies should expand the scope to include a broader range of protected attributes, LLM types (across different languages), and cultural contexts to understand bias patterns across various societal backgrounds.
Second, the prompts used in this study were consistent without variation or validation. Future studies could investigate the impact of prompt engineering on ethical decision-making, revealing how variations in tone, phrasing, and emotional content influence model responses.
Finally, this study used four-dimensional strategy to measure the bias in LLMs, which lack of the reflection and connection to reality. Future work could conduct comparative studies between human and LLM ethical decision-making to reveal the moral alignment between LLMs and human, providing insights for advancing human-AI interaction in ethical contexts.